The Northwood School Math Club is having a tamale sale. The school has 1600 students, but the club members are not sure how many tamales to make. One day during lunch, the club asked random students if they would buy a tamale. They found that 15 out of 80 students surveyed said they would definitely buy a tamale. How many tamales should the Math Club expect to sell? Use the Fraction Busters Method to solve the proportion:

Found 2 solutions by Boreal, Theo:
Answer by Boreal(15235)


15/80=x/1600
1600 is 20 times 80
Therefore, x is 20 times 15, or 300.
Alternatively, cross-multiply so that 80x=24000 and x=300

Answer by Theo(13342)

the school has 1600 students.
15 out of 80 students would definitely buy a tamale.
equation is:

15/80 = x/1600

multiply both sides of this equation by 1600 to get:

1600 * 15/80 = 1600 * x/1600

80 goes into 1600 20 times.
1600 goes into 1600 1 time.

equation becomes:

20 * 15 = x

solve for x to get x = 300.

300 students will most likely buy a tamale based on the survey.

check your fractions to see that they're true.

15 / 80 = 300 / 1600

cross multiply to get 15 * 1600 = 80 * 300

simplify to get 24000 = 24000

ratio is good.